OpenStudy (anonymous):

Suppose that G(x) = F(x - 6) - 3. Which statement best compares the graph of G(x) with the graph of F(x)? A. The graph of G(x) is the graph of F(x) shifted 6 units to the right and 3 units down. B. The graph of G(x) is the graph of F(x) shifted 6 units to the left and 3 units down. C. The graph of G(x) is the graph of F(x) shifted 6 units to the left and 3 units up. D. The graph of G(x) is the graph of F(x) shifted 6 units to the right and 3 units up.

OpenStudy (anonymous):

@study100

OpenStudy (xapproachesinfinity):

graph of g(x)=f(x-6)-3 means it is shifted 6 units to the right and 3 units down

OpenStudy (xapproachesinfinity):

that would be A explanation let f and g be two function g(x)=f(x-u) shifted u units to the right g(x)=f(x+u) shifted u units to the left g(x)=f(x)+u shifted u units up g(x)=f(x)-u shifted u units down

OpenStudy (anonymous):

If H(x) = 4x2 - 16 were shifted 5 units to the right and 2 down, what would the new equation be?